# The Power of Types: Exploring the Impact of Type Checking on Neural Bug Detection in Dynamically Typed Languages thanks: Partially supported by the FRQNT-B2X project (file number: 319955), IT30340 Mitacs Accelerate, and the Wallenberg AI, Autonomous Systems and Software Program (WASP), Sweden

### v1

@@ IV-C2 Results for the Real dataset
CodeBert 28.00 25.23 32.26 35.58
GraphCodeBert 30.84 27.13 32.59 35.81
UnixCoder 31.85 27.98 34.56 37.69
GGNN 10.51 11.31 17.29 21.81
GREAT 9.92 10.35 26.69 30.49
@@
Figure 6: Precision and recall of the bug detectors (a) and type checkers (b) on Annotated annotated with types for the NBD and pipeline performance on the unannotated (raw) and annotated version. F-scores are omitted since both precision and recall improve consistently when programs are annotated. Numbers are in percentage.
@@
NBDs in the Pipeline configuration demonstrate higher recall than those in the NN setup. Additionally, for certain models (GGNN and GREAT), the precision also improves. This is likely caused by the fact that the type checker has better precision than these NBDs. The recall improvement ranges from 3.13% to 4.52% for all NBDs, indicating that integrating a type checker can help identify more bugs. However, the precision in the Pipeline setting also decreases by between $2.77\%$ to $3.87\%$ for all pre-trained models due to the new false positives introduced by the type checker.
To better understand the trade-off between precision and recall introduced by the type checker on unannotated programs, we plot the F-$\beta$ score for different values of $\beta$ in 5(b). The two full training methods are excluded as the Pipeline approach improves both precision and recall. In general, when $\beta=1$, the NN setting achieves a higher score compared to the Pipeline setting, indicating that it may not be necessary to include type checking when the frequency of false alarms is a concern. However, when $\beta\geq 1.34$ (i.e., recall is considered $1.34$ times more important than precision), Pipeline surpasses the NN setting for all NBDs, suggesting that the type checker can be beneficial when recall of bugs is more important.
@@ IV-C3 Results for the Annotated dataset
Figure 6 displays the precision and recall of the bug detectors for both the NN and Pipeline settings, as well as the performance of both type checkers on the Annotated dataset. Additionally, we include the performance of these settings on the same dataset without type annotations (raw). This comparison is provided because the figures reported for the Real dataset in 5(a) are not directly comparable due to the Real dataset being highly imbalanced, whereas the Annotated dataset is balanced.
@@
RQ2: Combining a type checker with NBDs can improve recall by $3.13\%$ to $4.52\%$, but may lower precision by $2.77\%$ to $3.87\%$ on unannotated programs. On annotated programs, integrating a type checker improves both precision and recall. Thus, the pipeline setting is beneficial when recall is key or when type annotations are present.
@@ IV-D1 Addressing the RQ
TABLE II: Impact of filtering type-related bugs on precision and recall. Numbers are in percentage. F-scores are omitted since both precision and recall decrease consistently.
@@
CodeBert 93.40 91.91 -1.59 91.67 90.77 -0.98 28.00 26.96 -3.70 32.26 31.85 -1.29 74.49 69.51 -6.68 39.25 36.54 -6.90
GraphCodeBert 94.00 92.63 -1.46 92.19 91.38 -0.88 30.84 29.63 -3.94 32.59 32.12 -1.43 72.00 66.67 -7.41 39.78 37.09 -6.77
UnixCoder 94.07 92.70 -1.45 93.58 92.79 -0.84 31.85 30.68 -3.65 34.56 34.10 -1.31 76.24 73.81 -3.18 41.62 39.49 -5.12
GGNN 86.38 83.12 -3.77 80.89 77.31 -4.42 10.51 9.82 -6.56 17.29 16.86 -2.50 46.58 41.38 -11.16 19.88 16.44 -17.32
GREAT 84.80 81.44 -3.97 89.68 87.53 -2.39 9.92 9.17 -7.60 26.69 25.70 -3.71 39.60 34.52 -12.83 24.24 20.57 -15.16
@@ IV-D2 Results
Table II shows the precision and recall of different NBDs. Comparing the performance of the NBDs on synthetic and real-world datasets, we observe that these detectors perform much better on the synthetic datasets compared to the real-world bugs, consistent with previous work [41, 42]. This discrepancy can be attributed to the distribution shift between synthetic and real-world bugs [41].
@@
For unannotated programs, removing type-related bugs from the evaluation dataset leads to an average reduction in precision ($3.77\%$) and recall ($1.98\%$). The NBDs exhibit a more notable decrease in compared to Synthetic 1, with an average of $5.09\%$ and $2.05\%$ for precision and recall, respectively. With type annotations (Annotated), the decrease is even more notable, with the precision of NBDs dropping by $3.18\%$ to $12.83\%$ and recall dropping by $5.12\%$ to $17.32\%$.
@@
RQ3: Our findings reveal that bug detectors exhibit superior performance in detecting type-related bugs compared to other types of bugs. Removing these bugs from the evaluation dataset leads to a reduction in both precision (an average of 3.77%) and recall (an average of 1.98%) even in the absence of type annotations. When type annotation exists, the dropping in performance is even more significant. This suggests that the inclusion of type-related bugs in the evaluation dataset, particularly when using a type checker, may yield overly optimistic evaluation results.
@@ IV-E1 Addressing the RQ
TABLE III: Performance of detectors trained with full dataset $\rightarrow$ filtered dataset. Numbers are in percentage. In general, filtering type-related bugs in training improves the recall while decreasing the precision.
@@
CodeBert 91.91$\rightarrow$91.38 90.77$\rightarrow$91.17 91.33$\rightarrow$91.28 91.12$\rightarrow$91.23 26.96$\rightarrow$25.48 31.85$\rightarrow$31.67 29.20$\rightarrow$28.24 30.17$\rightarrow$29.47
GraphCodeBert 92.63$\rightarrow$92.12 91.38$\rightarrow$91.65 92.00$\rightarrow$91.88 91.76$\rightarrow$91.79 29.63$\rightarrow$29.25 32.12$\rightarrow$34.87 30.82$\rightarrow$31.82 31.31$\rightarrow$32.93
UnixCoder 92.70$\rightarrow$92.22 92.79$\rightarrow$93.10 92.75$\rightarrow$92.66 92.76$\rightarrow$92.83 30.68$\rightarrow$29.49 34.10$\rightarrow$36.60 32.30$\rightarrow$32.67 32.97$\rightarrow$34.08
GGNN 83.12$\rightarrow$82.78 77.31$\rightarrow$78.18 80.11$\rightarrow$80.42 79.01$\rightarrow$79.54 9.82$\rightarrow$9.42 16.86$\rightarrow$17.17 12.41$\rightarrow$12.17 13.82$\rightarrow$13.70
GREAT 81.44$\rightarrow$82.20 87.53$\rightarrow$87.72 84.38$\rightarrow$84.87 85.56$\rightarrow$85.95 9.17$\rightarrow$10.13 25.70$\rightarrow$27.19 13.51$\rightarrow$14.76 16.53$\rightarrow$17.91
@@ IV-E2 Results
The performance change of NBDs trained with the full dataset and the filtered dataset is shown in Table III. Values on the left side of the arrow show the performance of the full training dataset while the right side represents the performance of filtered training. The better F-$\beta$ scores of the two settings are highlighted in bold.
In general, the recall is improved when the NBD is trained on the filtered dataset (observed in 9/10 cases) with an average improvement of $1.01\%$. This improvement indicates that the removal of type-related bugs from the training data enables the NBDs to focus more on other non-type-related bugs. On the other hand, the precision is also reduced in 8 out of 10 cases. However, the change is generally smaller, with an average decrease of $0.66\%$. We hypothesize that this decrease may be attributed to the oversampling strategy we used, rather than generating entirely new synthetic samples.
When assessing the F-$\beta$ score, we evaluate both cases of $\beta=1$ (F-1) and $\beta=1.5$. When using the F-1 score, where recall is equally important as precision, filtering out type-related bugs from the training data does not yield a visible improvement in overall performance, as indicated by a better F-1 score in only 5 out of 10 cases. This finding aligns with our observation in Section IV-C that integrating type checkers is not beneficial in these cases.
When recall is more important ($\beta=1.5$), we note that the filtered dataset consistently outperforms the full dataset in all cases for the synthetic dataset. This observation suggests that the reduction in precision is less pronounced compared to the increase in recall in this case. This improvement extends to the real-world dataset as well, with the filtered dataset outperforming the full dataset in 3 out of 5 cases, while performing similarly in the case of GGNN. The slight variation in results for the real-world dataset may stem from the distribution shift between synthetic and real-world bugs.
RQ4: Removing type-related errors from the training dataset causes a significant increase in recall. Although a decrease in precision is also noted, the overall F-$\beta$ score is improved in 8 out of 10 cases when $\beta=1.5$.

### v2

@@ IV-C2 Results for the Real dataset
CodeBERT 27.91 24.54 32.29 34.83
GraphCodeBERT 30.80 26.48 32.71 35.02
UniXcoder 31.71 27.38 34.54 36.96
GGNN 10.51 11.02 17.47 21.44
GREAT 9.86 10.05 26.59 29.83
@@
TABLE II: Precision and recall of the bug detectors (a) and type checkers (b) on Annotated annotated with types for the NBD and pipeline performance on the unannotated (raw) and annotated version. F-scores are omitted since both precision and recall improve consistently when programs are annotated. Numbers are in percentage.
@@
NBDs in the Pipeline configuration demonstrate higher recall than those in the NN setup. Additionally, for certain models (GGNN and GREAT), the precision also improves. This is likely caused by the fact that the type checker has better precision than these NBDs. The recall improvement ranges from $2.31\%$ to $3.97\%$ for all NBDs, indicating that integrating a type checker can help identify more bugs. However, the precision in the Pipeline setting also decreases by between $3.37\%$ to $4.33\%$ for all pre-trained models due to the new false positives introduced by the type checker.
To better understand the trade-off between precision and recall introduced by the type checker on unannotated programs, we plot the F-$\beta$ score for different values of $\beta$ in 5(b). The two full training methods are excluded as the Pipeline approach improves both precision and recall. In general, when $\beta=1$, the NN setting achieves a higher score compared to the Pipeline setting, indicating that it may not be necessary to include type checking when the frequency of false alarms is a concern. However, when $\beta\geq 1.62$ (i.e., recall is considered $1.62$ times more important than precision), Pipeline surpasses the NN setting for all NBDs, suggesting that the type checker can be beneficial when recall of bugs is more important.
@@ IV-C3 Results for the Annotated dataset
5(d) displays the precision and recall of the bug detectors for both the NN and Pipeline settings, as well as the performance of both type checkers on the Annotated dataset. Additionally, we include the performance of these settings on the same dataset without type annotations (raw). This comparison is provided because the figures reported for the Real dataset in 5(a) are not directly comparable due to the Real dataset being highly imbalanced, whereas the Annotated dataset is balanced.
@@
RQ2: Combining a type checker with NBDs can improve recall by 2.31% to 3.97%, but may lower precision by $3.37\%$ to $4.33\%$ on unannotated programs. On annotated programs, integrating a type checker improves both precision and recall. Thus, the pipeline setting is beneficial when recall is key or when type annotations are present.
@@ IV-D1 Addressing the RQ
TABLE III: Impact of filtering type-related bugs on precision and recall. Numbers are in percentage. F-scores are omitted since both precision and recall decrease consistently.
@@
CodeBERT 93.40 91.91 -1.59 91.67 90.77 -0.98 27.91 26.10 -6.48 32.29 31.02 -3.95 74.49 69.51 -6.68 39.25 36.54 -6.90
GraphCodeBERT 94.00 92.63 -1.46 92.19 91.38 -0.88 30.80 28.80 -6.49 32.71 31.27 -4.42 72.00 66.67 -7.41 39.78 37.09 -6.77
UniXcoder 94.07 92.70 -1.45 93.58 92.79 -0.84 31.71 29.92 -5.67 34.54 33.32 -3.55 76.24 73.81 -3.18 41.62 39.49 -5.12
GGNN 86.38 83.12 -3.77 80.89 77.31 -4.42 10.51 9.47 -9.88 17.47 16.43 -5.93 46.58 41.38 -11.16 19.88 16.44 -17.32
GREAT 84.80 81.44 -3.97 89.68 87.53 -2.39 9.86 8.84 -10.35 26.59 24.95 -6.16 39.60 34.52 -12.83 24.24 20.57 -15.16
@@ IV-D2 Results
Table III shows the precision and recall of different NBDs. Comparing the performance of the NBDs on synthetic and real-world datasets, we observe that these detectors perform much better on the synthetic datasets compared to the real-world bugs, consistent with previous work [41, 42]. This discrepancy can be attributed to the distribution shift between synthetic and real-world bugs [41].
@@
For unannotated programs, removing type-related bugs from the evaluation dataset leads to an average reduction in precision ($5.11\%$) and recall ($3.35\%$). The NBDs exhibit a more notable decrease in compared to Synthetic 1, with an average of $7.77\%$ and $4.80\%$ for precision and recall, respectively. With type annotations (Annotated), the decrease is even more notable, with the precision of NBDs dropping by $3.18\%$ to $12.83\%$ and recall dropping by $5.12\%$ to $17.32\%$.
@@
RQ3: Our findings reveal that bug detectors exhibit superior performance in detecting type-related bugs compared to other types of bugs. Removing these bugs from the evaluation dataset leads to a reduction in both precision (an average of 5.11%) and recall (an average of 3.35%) even in the absence of type annotations. When type annotation exists, the dropping in performance is even more significant. This suggests that the inclusion of type-related bugs in the evaluation dataset, particularly when using a type checker, may yield overly optimistic evaluation results.
@@ IV-E1 Addressing the RQ
TABLE IV: Performance of detectors trained with full dataset $\rightarrow$ filtered dataset. Numbers are in percentage. In general, filtering type-related bugs in training improves the recall while decreasing the precision.
@@
CodeBERT 91.91$\rightarrow$91.38 90.77$\rightarrow$91.17 91.33$\rightarrow$91.28 91.12$\rightarrow$91.23 26.10$\rightarrow$24.77 31.02$\rightarrow$31.06 28.35$\rightarrow$27.56 29.32$\rightarrow$28.81
GraphCodeBERT 92.63$\rightarrow$92.12 91.38$\rightarrow$91.65 92.00$\rightarrow$91.88 91.76$\rightarrow$91.79 28.80$\rightarrow$28.69 31.27$\rightarrow$34.35 29.99$\rightarrow$31.26 30.47$\rightarrow$32.38
UniXcoder 92.70$\rightarrow$92.22 92.79$\rightarrow$93.10 92.75$\rightarrow$92.66 92.76$\rightarrow$92.83 29.92$\rightarrow$28.89 33.32$\rightarrow$35.98 31.52$\rightarrow$32.05 32.19$\rightarrow$33.45
GGNN 83.12$\rightarrow$82.78 77.31$\rightarrow$78.18 80.11$\rightarrow$80.42 79.01$\rightarrow$79.54 9.47$\rightarrow$9.33 16.43$\rightarrow$17.04 12.02$\rightarrow$12.06 13.40$\rightarrow$13.58
GREAT 81.44$\rightarrow$82.20 87.53$\rightarrow$87.72 84.38$\rightarrow$84.87 85.56$\rightarrow$85.95 8.84$\rightarrow$10.06 24.95$\rightarrow$27.12 13.05$\rightarrow$14.67 15.98$\rightarrow$17.82
@@ IV-E2 Results
The performance change of NBDs trained with the full dataset and the filtered dataset is shown in Table IV. Values on the left side of the arrow show the performance of the full training dataset while the right side represents the performance of filtered training. The better F-$\beta$ scores of the two settings are highlighted in bold.
In general, the recall is improved when the NBD is trained on the filtered dataset (observed in all cases) with an average improvement of $1.06\%$. This improvement indicates that the removal of type-related bugs from the training data enables the NBDs to focus more on other non-type-related bugs. On the other hand, the precision is also reduced in 8 out of 10 cases. However, the change is generally smaller, with an average decrease of $0.56\%$. We hypothesize that this decrease may be attributed to the oversampling strategy we used, rather than generating entirely new synthetic samples.
When assessing the F-$\beta$ score, we evaluate both cases of $\beta=1$ (F-1) and $\beta=1.5$. When using the F-1 score, where recall is equally important as precision, filtering out type-related bugs from the training data does not yield a visible improvement in overall performance, as indicated by a better F-1 score in only 6 out of 10 cases. This finding aligns with our observation in Section IV-C that integrating type checkers may not be beneficial in these cases.
When recall is more important ($\beta=1.5$), we note that the filtered dataset consistently outperforms the full dataset in all cases for the synthetic dataset. This observation suggests that the reduction in precision is less pronounced compared to the increase in recall in this case. This improvement extends to the real-world dataset as well, with the filtered dataset outperforming the full dataset in 4 out of 5 cases while not outperforming in the case of CodeBERT. The slight variation in results for the real-world dataset may stem from the distribution shift between synthetic and real-world bugs.
RQ4: Removing type-related errors from the training dataset causes a significant increase in recall. Although a decrease in precision is also noted, the overall F-$\beta$ score is improved in 9 out of 10 cases when $\beta=1.5$.